zepdrix (zepdrix):

PDE - Change of coordinate axes. I'm confused how these second partials are calculated. Maybe someone can me remind me. I think it's from Calc3 stuff. \[\Large \xi = x\cos\alpha+y\sin\alpha\]\[\Large \eta = -x\sin\alpha+y\cos\alpha\] \[\Large u(x,y)\to u\left[\xi(x,y),\eta(x,y)\right]\] First partials:\[\Large u_x = u_{\xi}\xi_x+u_{\eta}\eta_x\]\[\Large u_y = u_{\xi}\xi_y+u_{\eta}\eta_y\] Which simplify to:\[\Large u_x = \cos\alpha u_{\xi}-\sin\alpha u_{\eta}\]\[\Large u_y = \sin\alpha u_{\xi}+\cos\alpha u_{\eta}\]

zepdrix (zepdrix):

Teacher wrote that the second partials are giving us ...\[\Large u_{xx}= \cos^2\alpha (u_{\xi\xi})\color{royalblue}{-2\sin \alpha \cos \alpha (u_{\xi \eta})}+\sin^2\alpha(u_{\eta \eta})\]

zepdrix (zepdrix):

I guess I'm especially confused on that middle term.. Maybe I'm just not remembering how to chain rule these functions of multiple variables.

zepdrix (zepdrix):

\[\Large u_{yy}=\sin^2\alpha (u_{\xi \xi})+2\sin \alpha \cos \alpha (u_{\xi \eta})+\cos^2\alpha (u_{\eta \eta})\]

zepdrix (zepdrix):

Phew! Luckily it wasn't on the test. :O